7. A firm pays a $12.80 dividend at the end of year one (D1), has a...

7.

A firm pays a $12.80 dividend at the end of year one (D1), has a stock price of $88, and a constant growth rate (g) of 5 percent.
  
Compute the required rate of return (Ke). (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)
  

8.

The Pioneer Petroleum Corporation has a bond outstanding with an $100 annual interest payment, a market price of $890, and a maturity date in five years. Assume the par value of the bond is $1,000.  
    
Find the following: (Use the approximation formula to compute the approximate yield to maturity and use the calculator method to compute the exact yield to maturity. Do not round intermediate calculations. Input your answers as a percent rounded to 2 decimal places.)
Answer #1

Answer for question 7

Required rate of return = [Dividend / stock price] + growth rate

Dividend = 12.80

Stock price = 88

Growth rate = 0.05

Required rate of return = [12.80/88]+0.05 = 0.1454+0.05 = 0.1954 = 19.54%

Required rate of return =19.54%